OpenStudy (anonymous):

Determine if Rolle's Theorem can be applied to the function on the given interval; if so, find the values of c guaranteed by the theorem. f(x)= x^2/3 - 1 on [-8,8]

OpenStudy (jennychan12):

continuous and differential on [-8,8] and f(-8)=f(-8) then take f'(x) = 0.

OpenStudy (anonymous):

don't forget to do the first part. that is, don't forget to see if \(f\) is differentiable on \((-8,8)\) which means the derivative has to exist for all numbers in that interval take the derivative, and check hint: it isn't

OpenStudy (anonymous):

I got 2/3x^1/3 = 3...?

OpenStudy (anonymous):

exponent should be negative

OpenStudy (anonymous):

what do you next to solve?

OpenStudy (jennychan12):

i don't think there is an answer cuz then 2/3x^1/3 =0 has no solution.

OpenStudy (anonymous):

there is no answer because it doesn't satisfy the hypotheses the derivative does not exist at \(x=0\)

OpenStudy (anonymous):

the derivative is \[\frac{2}{3\sqrt[3]{x}}\] and this is undefined at \(x=0\) so the function is not differentiable at zero

OpenStudy (anonymous):

and therefore it violates the hypothesis of rolles theorem, which requires the function to be differentiable on the entire interval

OpenStudy (anonymous):

thank you!(: